# Conformalized Bayesian Inference, with Applications to Random Partition Models

### v1

@@ 4 Real data applications
We are now in a position to illustrate our CBI methodology, in its $\mathcal{D}_{VI}$-based instantiation for random partition models, as applied to three real-world datasets: the classic Galaxy velocities data (Roeder,, 1990), colorectal cancer spatial transcriptomics data (Lee et al.,, 2020; Duan et al.,, 2025), and ERP functional data (Song et al.,, 2025; Kappenman et al.,, 2021).
@@ 4.1 Galaxy velocities data
The Galaxy velocities dataset is a well-known and simple benchmark for mixture-based density estimation and clustering. It contains velocity measurements (in units of 1,000 km/second) for 82 galaxies observed in six well-separated conic sections of the Corona Borealis region (Roeder,, 1990). We analyze this dataset using a Pitman–Yor Gaussian mixture model, fitted via a marginal Gibbs sampler implemented in pyrichlet (Selva et al.,, 2025). For the subsequent CBI analysis, we use 5,000 training and 1,000 calibration MCMC samples after appropriate burn-in and thinning. Figure 3(a) shows a histogram of the data together with the posterior predictive estimate of the density.
Figure 3(b) visualizes the estimated posterior distribution of the number of clusters in the data, which is spread over a relatively wide range of values, indicating substantial uncertainty. Figure 3(c) displays the $\mathcal{D}_{VI}$-KDE point estimate of the data partition, which reveals three spatially well-separated clusters. At first glance, this may appear inconsistent with the relatively small posterior mass assigned to partitions with three clusters in Figure 3(b). However, recall that our point estimation procedure seeks the partition maximizing the $\mathcal{D}_{VI}$-KDE score. Since this score is constructed using the VI distance, the procedure naturally selects a representative partition that is well supported by the posterior training samples under that distance. Because, as we have already discussed, the very design of $\mathcal{D}_{VI}$ favors more parsimonious partitions, this result is in fact expected and perhaps even desirable: in the absence of strong prior beliefs supporting the presence of additional clusters, an Occam’s razor principle naturally suggests preferring a simpler clustering representation. This is particularly reasonable if many of the sampled partitions, for example those with four or five clusters, differ only slightly from the selected three-cluster configuration (see also the point coloring in Figure 4(a)).
@@
(c) $\mathcal{D}_{VI}$-KDE point estimator of the data clustering (coloring by cluster membership).
The view that the posterior distribution reflects a substantial degree of uncertainty is further supported by the $90\%$ conformal credible set, which we find to include all $k$-means clusterings with $k=3,4,5,6$. Additionally, we repeat the size check performed in the simulated data example of the previous section, where we test whether 1,000 randomly generated partitions fall within the conformal set. The results are encouraging, as none of the randomly generated partitions fall inside the conformal set.
Finally, Figures 4(a) and 4(b) report the results from the DPC-based multimodality analysis. As the decision graph clearly shows, two well-separated modes are identified, corresponding to the three-cluster point estimate as well as a six-cluster refinement of it. Both partitions are plausible given the empirical distribution of the data, and the multimodality analysis reveals that both receive support by the posterior distribution over data partitions derived from the adopted Bayesian model.
@@
4.2 Colorectal cancer spatial transcriptomics data
In a recent study, Duan et al., (2025) analyze spatial transcriptomics data on colorectal cancer (Lee et al.,, 2020), with the aim of discovering spatially aligned subpopulations of tumor, immune, and stromal cells. Spatial alignment refers to the spatial co-localization of clusters across cell types, where spatial coordinates correspond to the two-dimensional positions of cells within a medical image. To capture this structure, the authors propose a spatially-aligned random partition model that combines a PY Gaussian mixture model for gene expression features with a prior dependence structure that induces spatial co-localization in the latent clustering across cell types. Their MCMC output consists of sampled partitions for each cell type: cells of different types are clustered separately, but the partitions are modeled as dependent through prior-induced spatial alignment. The results of their analysis indeed reveal subpopulations (clusters) of cells from different types that are spatially co-localized.
The left-hand side plots in Figure 6 show the $\mathcal{D}_{VI}$-KDE point estimates for tumor and immune cell partitions. Recall that the model of Duan et al., (2025) performs density estimation and clustering on high-dimensional gene expression features while using spatial coordinates to induce alignment. For visualization purposes, we display partitions only in the spatial domain, since spatial alignment is the focus of the analysis. Both point estimates reveal spatially coherent clusters, in line with the findings of Duan et al., (2025).
Finally, we use the conformal credible regions (constructed separately for both marginal posteriors) to test a hypothesis of global spatial alignment between the partitions of different cell types. Specifically, given the point estimate for type A cells (either tumor or immune), we spatially translate it to type B cells by matching each type B cell to its closest (spatially) type A cell, and assigning that type B cell to the cluster of the matched type A cell. The right-hand side plots in Figure 6 illustrate these translated partitions. We then check whether each translated partition of type B cells lies within the 90% conformal credible region for the posterior distribution over partitions of type B cells. We emphasize that this constitutes a test of global alignment, as it requires the entire translated partition to lie within the credible region. This is distinct from the goal of Duan et al., (2025), who adopt a more local notion of alignment seeking to identify subsets of clusters that are spatially co-located, rather than assuming full alignment across all clusters. As a result of our analysis, both spatially translated partitions fall outside the corresponding credible regions, supporting the hypothesis that only subsets of tumor and immune cell clusters are spatially co-located. This is consistent with the findings of Duan et al., (2025).
@@ 4.3 ERP data experiment
Our final application of CBI concerns functional data clustering using a specialized Bayesian random partition model. Specifically, we build on the analysis of Song et al., (2025), who employ a projection Determinantal Point Process (pDPP) repulsive mixture prior (Xu et al.,, 2016; Xie and Xu,, 2020; Petralia et al.,, 2012; Beraha et al.,, 2025; Cremaschi et al.,, 2025) to analyze data from the publicly available ERP CORE project (Kappenman et al.,, 2021). The goal of the analysis is to obtain interpretable clusters of functional data representing ERP waveforms recorded from 34 young adult participants exposed to certain visual stimuli. We refer the reader to Section 7 of Song et al., (2025) for additional scientific context. Figure 7(a) visualizes the analyzed data in terms of waveform amplitudes measured up to 800 milliseconds after stimulus onset. Song et al., (2025) fit their Bayesian pDPP repulsive mixture model using the first four functional principal components of the data. The employed MCMC algorithm generates posterior samples of the random partition of the 34 subjects, which we use as the input to our CBI analysis.
@@
As shown in Figure 7(b), the posterior distribution concentrates around a single representative partition, corresponding to the $\mathcal{D}_{VI}$-KDE point estimate in Figure 7(c), which features four distinct clusters. This estimate closely matches that obtained by Song et al., (2025) (see their Figure 5) using the SALSO algorithm (Dahl et al.,, 2022).
@@
After constructing a $90\%$ credible region from the calibration scores, we test two extreme hypotheses. First, we assess whether the posterior supports full heterogeneity, i.e., the trivial partition with 34 clusters. It does not: the corresponding conformal $p$-value [^8] Recall from Equation (6) that $\hat{p}(\theta;\bm{\theta}^{2})$ is a $p$-value under the null hypothesis $(\bm{\theta}^{2},\theta)\overset{\text{iid}}{\sim}\Pi$. equals $0.000999$, indicating that such a configuration is highly atypical under the posterior. This corroborates the finding by Song et al., (2025) that participant subpopulations exhibit meaningful homogeneity. Next, we test whether complete homogeneity (a single cluster) can be rejected as atypical under the posterior. Interestingly, the one-cluster partition is included in the credible set, with conformal $p$-value equal to $0.140859$, indicating that while it is not highly typical, it cannot be decisively ruled out. A robustness check using 1,000 partitions generated uniformly at random confirms that none are included in the credible region, reassuring of the test power. We nevertheless emphasize that this $p$-value is valid only under the specific null hypothesis that the one-cluster partition is a posterior draw independent of the calibration samples. Even ignoring this formal interpretation, the result should be strictly understood in terms of posterior typicality under the VI metric: the $\mathcal{D}_{VI}$-KDE score blends empirical posterior mass with the VI geometry, implying not that the one-cluster partition is likely (in fact, it never appears among the MCMC samples we use), but that it is not exceedingly atypical relative to posterior samples under the VI metric. More broadly, this null finding provides a good reminder that, while the KDE scoring rule yields a valid credible region with guaranteed posterior coverage by leveraging an informative metric such as $\mathcal{D}_{VI}$ to summarize a complex space like that of partitions, the results of conformal testing should be interpreted strictly within their intended scope.
@@
Figure 7: CBI results for ERP data analyzed using the pDPP mixture model of Song et al., (2025).
@@ 5 Discussion
This article has introduced Conformalized Bayesian Inference (CBI), a simple, robust and efficient pipeline to turn Monte Carlo samples from a Bayesian posterior distribution on a complex parameter space into interpretable insights about a single representative of that posterior, posterior credible regions based on conformal prediction ideas, and possible posterior multimodalities, all simply requiring a notion of discrepancy between sampled parameters. We have explored theoretical and methodological aspects of CBI, and illustrated its power on the notoriously difficult task of inferring the clustering structure of the data using Bayesian random partition models. We now turn to a discussion of the limitations of the methodology and what venues for future research are left open.
@@ 5.1 Current limitations
Second, although we have shown that the distance-based KDE score is a simple and robust choice for CBI, some posterior distributions may not be well captured by any readily available distance combined with a simple kernel. This limitation could be addressed by developing more flexible density estimators, for instance when the parameter space, such as the space of graphs, is suitable for modern deep learning architectures (Bishop and Bishop,, 2024).
Finally, we note that the current scope of CBI is limited to a single posterior distribution. However, much recent work (especially in Bayesian nonparametrics) has focused on dependent priors, and hence dependent posteriors, to model heterogeneous but related populations (MacEachern,, 2000; Teh et al.,, 2006; Rodriguez et al.,, 2008; Lijoi et al.,, 2014, among many others). For example, in our spatial transcriptomics experiment, the two marginal posteriors are dependent by construction, yet our analysis employed heuristic simplifications to treat each inference problem independently. A fully rigorous treatment of dependent posteriors remains an open problem and would substantially extend the applicability of CBI.
@@ 5.2 Future directions
Future research should further investigate the use of CBI on complex parameter spaces beyond partitions, such as graphs, matrices, regression functions, and mixing measures. As noted earlier, we provide preliminary empirical illustrations in the Supplementary Material for cases where $\Theta$ represents either the space of mixing measures in mixture models or the space of covariance matrices in covariance estimation problems. Additional methodological developments are also of interest, including the integration of alternative density-based clustering methods (Ester et al.,, 1996; Wang et al.,, 2019), which could offer a more refined characterization of posterior multimodality—distinguishing, for example, whether the identified representative parameters correspond to distinct high posterior-mass regions separated by low-mass areas, or instead belong to a single, spread-out region of posterior concentration.
Supplementary Material for “Conformalized Bayesian Inference, with Applications to Random Partition Models”
This appendix collects the supplementary material for the article “Conformalized Bayesian Inference, with Applications to Random Partition Models.” In particular, it reports the proofs of two propositions presented in the paper (Section A) as well as the results of further numerical experiments on Conformalized Bayesian Inference applied to (i) a unimodal posterior distribution on data partitions (Section B), (ii) a posterior on the space of mixing measures derived from a density mixture model (Section C), and (iii) a posterior on the space of Gaussian covariance matrices for multivariate data (Section D).
Appendix A Mathematical proofs
In this section, we present the proofs of Propositions 1 and 2.
A.1 Proof of Proposition 1
Equations (4) and (6) in the main text are classic results in conformal prediction. Hence, we omit their proof and refer the reader to classic references such as (Vovk et al.,, 2005; Angelopoulos and Bates,, 2023). Instead, we prove the concentration inequality in Equation (5) in the main text.
Step 1. The inclusion criterion for the credible region is equivalent to $s(\theta;\bm{\theta}^{1})\geq s_{(k)}$, where $s_{(k)}$ is the $k$-th order statistic of the calibration scores and $k=\lceil\alpha(N+1)-1\rceil$. Let $\Pi_{s}(s):=\Pi(\{\theta\in\Theta:s(\theta;\bm{\theta}^{1})\leq s\})$ denote the posterior CDF of the scores. The posterior mass of the random credible set is precisely
$\Pi(\mathcal{C}_{1-\alpha}(\bm{\theta}^{2}))=1-\Pi(\{\theta:s(\theta;\bm{\theta}^{1})<s_{(k)}\})=1-\Pi_{s}(s_{(k)}^{-}),$
where the minus sign denotes a left-hand limit. The quantity to be bounded is therefore $|\Pi(\mathcal{C}_{1-\alpha}(\bm{\theta}^{2}))-(1-\alpha)|=|\alpha-\Pi_{s}(s_{(k)}^{-})|$.
Step 2. For any $\delta\in(0,1)$, the DKW inequality (Dvoretzky et al.,, 1956; Massart,, 1990) gives
$\mathbb{P}\left(\sup_{s\in\mathbb{R}}|\hat{\Pi}_{s,N}(s)-\Pi_{s}(s)|\leq\epsilon_{N}(\delta)\right)\geq 1-\delta,$
@@ A.1 Proof of Proposition 1
    Empirical jump size: The term $|\frac{k}{N}-\hat{\Pi}_{s,N}(s_{(k)}^{-})|$ represents the proportion of samples with scores exactly equal to $s_{(k)}$. This is because $\hat{\Pi}_{s,N}(s_{(k)}^{-})=\frac{1}{N}|\{t:s(\theta_{t};\bm{\theta}^{1})<s_{(k)}\}|$. If $m$ samples have a score of $s_{(k)}$, then there are kmk-m scores strictly less than $s_{(k)}$, so $\hat{\Pi}_{s,N}(s_{(k)}^{-})=(k-m)/N$. The term thus equals $m/N$. In the absence of ties ($m=1$), this term is $1/N$, so it vanishes deterministically if the posterior distribution of the scores is continuous.
@@ A.2 Proof of Proposition 2
For any fixed history $\vec{\theta}_{<k}$, the inner term is an expectation difference bounded by the total variation distance between the corresponding measures:
@@
Appendix B Additional experiment 1: unimodal posterior distribution over partitions
We run a Bayesian clustering experiment analogous to the one used to illustrate CBI in the main text, using the same simulated two-dimensional dataset (Figure 8(a)). The only difference is that the Pitman-Yor (PY) Gaussian mixture prior is now specified with a smaller concentration parameter, set to $0.01$. As is well known (De Blasi et al.,, 2013), lowering this parameter increases prior mass on partitions with fewer clusters. This shift is reflected in the posterior over the number of clusters $K$, which becomes more concentrated on small values. Consistent with this, the $\mathcal{D}_{VI}$-KDE point estimate remains the same two-cluster configuration identified in the example with higher concentration parameter (Figures 8(b) and 8(c)).

### v2

@@ 6 Real data applications
We are now in a position to illustrate our CBI methodology, in its $\mathcal{D}_{VI}$-based instantiation for random partition models, as applied to two real-world datasets: the classic Galaxy velocities data (roeder1990density) and ERP functional data (song2025repulsive, kappenman2021erp).
@@ 6.1 Galaxy velocities data
The Galaxy velocities dataset is a well-known and simple benchmark for mixture-based density estimation and clustering. It contains velocity measurements (in units of 1,000 km/second) for 82 galaxies observed in six well-separated conic sections of the Corona Borealis region (roeder1990density). We analyze this dataset using a Pitman–Yor Gaussian mixture model, fitted via a marginal Gibbs sampler implemented in pyrichlet (selva2025pyrichlet). For the subsequent CBI analysis, we use 5,000 training and 1,000 calibration MCMC samples after appropriate burn-in and thinning. Figure 5(a) shows a histogram of the data together with the posterior predictive estimate of the density.
Figure 5(b) visualizes the estimated posterior distribution of the number of clusters in the data, which is spread over a relatively wide range of values, indicating substantial uncertainty. Figure 5(c) displays the $\mathcal{D}_{VI}$-KDE point estimate of the data partition, which reveals three spatially well-separated clusters. At first glance, this may appear inconsistent with the relatively small posterior mass assigned to partitions with three clusters in Figure 5(b). However, recall that our point estimation procedure seeks the partition maximizing the $\mathcal{D}_{VI}$-KDE score. Since this score is constructed using the VI distance, the procedure naturally selects a representative partition that is well supported by the posterior training samples under that distance. Because, as we have already discussed, the very design of $\mathcal{D}_{VI}$ favors more parsimonious partitions, this result is in fact expected and arguably desirable: in the absence of strong prior beliefs supporting the presence of additional clusters, an Occam’s razor principle naturally suggests preferring a simpler clustering representation. This is particularly reasonable if many of the sampled partitions, for example those with four or five clusters, differ only slightly from the selected three-cluster configuration (see also the point coloring in Figure 6(a)).
@@
(c) $\mathcal{D}_{VI}$-KDE point estimator (coloring by cluster membership).
The view that the posterior distribution reflects a substantial degree of uncertainty is further supported by the $90\%$ conformal credible set, which we find to include all $k$-means clusterings with $k=3,4,5,6$. Additionally, we repeat the size check performed for the simulated data example in the previous section, where we test whether 1,000 randomly generated partitions fall within the conformal set. The results are encouraging, as none of the randomly generated partitions fall inside the conformal set.
Finally, Figures 6(a) and 6(b) report the results from the DPC-based multimodality analysis. As the decision graph clearly shows, two well-separated modes are identified, corresponding to the three-cluster point estimate as well as a six-cluster refinement of it. Both partitions are plausible given the empirical distribution of the data, and the multimodality analysis reveals that both receive support by the posterior distribution over data partitions derived from the adopted Bayesian model.
@@ 6.2 ERP data experiment
Our second application of CBI concerns functional data clustering using a specialized Bayesian random partition model. Specifically, we build on the analysis of song2025repulsive, who employ a projection Determinantal Point Process (pDPP) repulsive mixture prior (xu2016bayesian, xie2020bayesian, petralia2012repulsive, beraha2025bayesian, cremaschi2025repulsion) to analyze data from the publicly available ERP CORE project (kappenman2021erp). The goal of the analysis is to obtain interpretable clusters of functional data representing ERP waveforms recorded from 34 young adult participants exposed to certain visual stimuli. We refer the reader to Section 7 of song2025repulsive for additional scientific context. Figure 7(a) visualizes the analyzed data in terms of waveform amplitudes measured up to 800 milliseconds after stimulus onset. song2025repulsive fit their Bayesian pDPP repulsive mixture model using the first four functional principal components of the data. The employed MCMC algorithm generates posterior samples of the random partition of the 34 subjects, which we use as the input to our CBI analysis.
@@
As shown in Figure 7(b), the posterior distribution concentrates around a single representative partition, which corresponds to to the $\mathcal{D}_{VI}$-KDE point estimate in Figure 7(c) featuring four distinct clusters. Interestingly, this estimate matches that obtained by song2025repulsive (see their Figure 5) using the SALSO algorithm (dahl2022search).
@@
After constructing the $90\%$ credible region from the calibration scores, we test two extreme hypotheses. First, we assess whether the posterior supports full heterogeneity, i.e., the trivial partition with 34 clusters. It does not: the corresponding conformal $p$-value [^13] Recall from Equation (6) that $\hat{p}(\theta;\boldsymbol{\theta}^{2})$ is a $p$-value under the null hypothesis $(\boldsymbol{\theta}^{2},\theta)\overset{\text{iid}}{\sim}\Pi$. equals $0.000999$, indicating that such a configuration is highly atypical under the posterior. This corroborates the finding by song2025repulsive that participant subpopulations exhibit meaningful homogeneity. Next, we test whether complete homogeneity (a single cluster) can be rejected as atypical under the posterior. Interestingly, the one-cluster partition is included in the credible set, with conformal $p$-value equal to $0.140859$, indicating that while it is not highly typical, it cannot be decisively ruled out. A robustness check using 1,000 partitions generated uniformly at random confirms that none are included in the credible region, reassuring of the test power. We nevertheless emphasize that this $p$-value is valid only under the specific null hypothesis that the one-cluster partition is a posterior draw independent of the calibration samples. Even ignoring this formal interpretation, the result should be strictly understood in terms of posterior typicality under the VI metric: the $\mathcal{D}_{VI}$-KDE score blends empirical posterior mass with the VI geometry, implying not that the one-cluster partition is likely (in fact, it never appears among the MCMC samples we use), but that it is not exceedingly atypical relative to posterior samples under the VI metric. More broadly, this null finding provides a good reminder that, while the KDE scoring rule yields a valid credible region with guaranteed posterior coverage by leveraging an informative metric such as $\mathcal{D}_{VI}$ to summarize a complex space like that of partitions, the results of conformal testing should be interpreted strictly within their intended scope.
@@
Figure 7: CBI results for ERP data analyzed using the pDPP mixture model of song2025repulsive.
@@ 7 Discussion
This article has introduced Conformalized Bayesian Inference (CBI), a simple, robust and efficient pipeline to turn Monte Carlo samples from a Bayesian posterior distribution on a complex parameter space into interpretable insights about a single representative of that posterior, a posterior credible region based on conformal prediction ideas, and possible posterior multimodalities, all simply requiring a notion of discrepancy between sampled parameters. We have explored theoretical and methodological aspects of CBI, and illustrated its power on the notoriously difficult task of inferring the clustering structure of data using Bayesian random partition models. We now turn to a discussion of the limitations of the methodology and what venues for future research are left open.
@@ 7.1 Current limitations
Second, although we have shown that the distance-based KDE score is a simple and robust choice for CBI, some posterior distributions may not be well captured by any readily available distance combined with a simple kernel. This limitation could be addressed by developing more flexible pseudo-density estimators, for instance when the parameter space (such as the space of graphs) is suitable for modern deep learning architectures (bishop2024deeplearning).
Finally, we note that the current scope of CBI is limited to a single posterior distribution. However, much recent work (especially in Bayesian nonparametrics) has focused on dependent priors, and hence dependent posteriors, to model heterogeneous but related populations (maceachern2000dependent, teh2006hierarchical, rodriguez2008nested, lijoi2014bayesian, among many others). A rigorous treatment of dependent posteriors remains an open problem and would substantially extend the applicability of CBI.
@@ 7.2 Future directions
Future research should further investigate the use of CBI on complex parameter spaces beyond partitions, such as graphs, matrices, regression functions, and mixing measures. As noted earlier, we provide preliminary empirical illustrations in the Supplementary Material for cases where $\Theta$ represents either the space of mixing measures in mixture models (where $\mathcal{D}$ coincides with the Wasserstein-1 metric), or the space of covariance matrices (discrepancies being measured according to the operator norm) in covariance estimation problems. Additional methodological developments are also of interest, including the integration of alternative density-based clustering methods (ester1996density, wang2019dbscan), which could offer a more refined characterization of posterior multimodality—distinguishing, for example, whether the identified representative parameters correspond to distinct high posterior-mass regions separated by low-mass areas, or instead belong to a single, spread-out region of posterior concentration.
SUPPLEMENTARY MATERIAL
7.3 Proof of Proposition 1
Equations (4) and (6) in the main text are classic results in conformal prediction. Hence, we omit their proof and refer the reader to classic references such as vovk2005algorithmic, angelopoulos2023conformal. Instead, we prove the concentration inequality in Equation (5).
Step 1. The inclusion criterion for the credible region is equivalent to $s(\theta;\boldsymbol{\theta}^{1})\geq s_{(k)}$, where $s_{(k)}$ is the $k$-th order statistic of the calibration scores and $k=\lceil\alpha(N+1)-1\rceil$. Let $\Pi_{s}(s):=\Pi(\{\theta\in\Theta:s(\theta;\boldsymbol{\theta}^{1})\leq s\})$ denote the posterior CDF of the scores. The posterior mass of the random credible set is precisely
$\Pi(\mathcal{C}_{1-\alpha}(\boldsymbol{\theta}^{2}))=1-\Pi(\{\theta:s(\theta;\boldsymbol{\theta}^{1})<s_{(k)}\})=1-\Pi_{s}(s_{(k)}^{-}),$
where the minus sign denotes a left-hand limit. The quantity to be bounded is therefore $|\Pi(\mathcal{C}_{1-\alpha}(\boldsymbol{\theta}^{2}))-(1-\alpha)|=|\alpha-\Pi_{s}(s_{(k)}^{-})|$.
Step 2. For any $\delta\in(0,1)$, the DKW inequality (dvoretzky1956asymptotic, massart1990tight) gives
$\mathbb{P}_{\boldsymbol{\theta}^{2}\,\overset{\textnormal{iid}}{\sim}\,\Pi}\left(\sup_{s\in\mathbb{R}}|\hat{\Pi}_{s,N}(s)-\Pi_{s}(s)|\leq\epsilon_{N}(\delta)\right)\geq 1-\delta,$
@@ 7.3 Proof of Proposition 1
    Empirical jump size: The term $|\frac{k}{N}-\hat{\Pi}_{s,N}(s_{(k)}^{-})|$ represents the proportion of samples with scores exactly equal to $s_{(k)}$. This is because $\hat{\Pi}_{s,N}(s_{(k)}^{-})=\frac{1}{N}|\{t:s(\theta_{t};\boldsymbol{\theta}^{1})<s_{(k)}\}|$. If $m$ samples have a score of $s_{(k)}$, then there are kmk-m scores strictly less than $s_{(k)}$, so $\hat{\Pi}_{s,N}(s_{(k)}^{-})=(k-m)/N$. The term thus equals $m/N$. In the absence of ties ($m=1$), this term is $1/N$, so it vanishes deterministically if the posterior distribution of the scores is continuous.
@@ 7.4 Proof of Proposition 2
For any fixed history $\vec{\theta}_{<k}$, the inner term is a difference of expectations bounded by the TV distance between the corresponding measures:
@@
8 Additional experiment 1: unimodal posterior distribution over partitions
We run a Bayesian clustering experiment analogous to the one used to illustrate CBI in the main text, using the same simulated two-dimensional dataset (Figure 8(a)). The only difference is that the Pitman-Yor (PY) Gaussian mixture prior is now specified with a smaller concentration parameter, set to $0.01$. As is well known (deblasi2013gibbs), lowering this parameter increases prior mass on partitions with fewer clusters. This shift is reflected in the posterior over the number of clusters $K$, which becomes more concentrated on small values. Consistent with this, the $\mathcal{D}_{VI}$-KDE point estimate remains the same two-cluster configuration identified in the example with higher concentration parameter (Figures 8(b) and 8(c)).